A ball is thrown straight upward with an initial velocity of 28 m/s and an initial height of 35 m. how much time does it take for the ball to land
1
Expert's answer
2018-09-06T16:51:08-0400
H1 = 35m
V0 = 28m/s
t1 = (V-V_0)/(-g)=2.8s
S1 = (V^2-V_0^2)/(-2g)=39.2m
S = H1 + S1 = 74.2m
t = t_1+t_2=t_1+ √2S/√g=6.652s